OpenStudy (nicole143):

@mathmale How do I write this in simplest form? Square root 2 - square root 6 / square root 2 + square root 6

OpenStudy (nicole143):

\[\sqrt{2} - \sqrt{6} / \sqrt{2} + \sqrt{6} \]

OpenStudy (mathmale):

I'd be very interested in seeing what you've done already.

OpenStudy (mathmale):

Great minds think alike, don't they? :)

OpenStudy (nicole143):

All I did was take out the individual square roots and swap them for a big one over the whole thing.

OpenStudy (mathmale):

Excuse my pickiness, but you do have to enclose the first 2 and the last 2 terms inside parentheses. See why?

OpenStudy (nicole143):

I don't see why

OpenStudy (mathmale):

This\[\frac{ \sqrt{2}-\sqrt{6} }{ \sqrt{2}+\sqrt{6} }\] is my interpretation of the original problem:

OpenStudy (mathmale):

Is it your interpretation also?

OpenStudy (nicole143):

Yes(:

OpenStudy (mathmale):

Forgive my temporary absence. My point is that if you don't enclose the first 2 terms and then the last 2 terms in ( ), some readers may not grasp that we intended for the first 2 terms to constitute one quantity and the last 2 terms another quantitiy.

OpenStudy (nicole143):

Oh, I see what you mean.

OpenStudy (mathmale):

In this particular problem, Nicole, you want to eliminate the radicals fromthe denominator. That's what "simplest form" means here. The conjugate of Sqrt(2) + Sqrt(6) is Sqrt( 2) - Sqrt (6). Multiply top and bottom of your fraction by that. You'll end up with an interger quantity in the denom, which is what we wanted.

OpenStudy (mathmale):

Lo and behold, no more radicals will exist in the denom.

OpenStudy (nicole143):

Okay, one moment

OpenStudy (nicole143):

Will you still have the roots in top? Or do I times both top and bottom?

OpenStudy (mathmale):

Yes, you'll still have some radicals in the numerator, which is fine. Radicals in the denom. are not so fine. ;)

OpenStudy (nicole143):

Okay ha

OpenStudy (mathmale):

I never crack jokes, you know. ;)

OpenStudy (nicole143):

Never? :P So because it Sqrt(2) + Sqrt(6) times Sqrt( 2) - Sqrt (6) it will coming out to be 2 - 6 on bottom?

OpenStudy (mathmale):

Exactly, which reduces to what negative number?

OpenStudy (nicole143):

-4

OpenStudy (mathmale):

Oh, really? Yes, of course, you're right.

OpenStudy (nicole143):

So after you get -4 on the bottom, you have to simplify the top. do I take it out the roots?

OpenStudy (mathmale):

No. combine similar terms as far as possible, and then leave the resulting fraction as is. Or, get rid of that (-) sign in the denom and put it in the num. instead.

OpenStudy (nicole143):

Can you show the steps, if any, after w get -4?

OpenStudy (nicole143):

*we

OpenStudy (mathmale):

Yeah, we. We make a good team. You already have a numerator, I assume. Simply enclose it in parentheses and stick a (-) sign in front of it, rewriting the den. as simply 4. You could leave your result like that, or you could distribute that (-) sign over to everything within your set of parenthteses.

OpenStudy (nicole143):

Haha:P I unfortunately don't have that as an answer choice:/ Unless i'm reading it wrong and the image of what it looks like isn't clicking in my head..

OpenStudy (nicole143):

Maybe I did something wrong? @mathmale

OpenStudy (mathmale):

I haven't actually done the problem myself, Nicole, since we were certainly on the right track. Is it possible for you to share an image of what you've written?

OpenStudy (nicole143):

Yes, one second.

OpenStudy (mathmale):

Seeing this image is SO helpful. Nicole, you need to multiply BOTH numerator and denom. of your fraction by Sqrt(2)-Sqrt(6). Mind trying again?

OpenStudy (nicole143):

oh, ha yes i redo it :P

OpenStudy (nicole143):

*I'll

OpenStudy (nicole143):

\[\sqrt{4} + \sqrt{36} \] For the top?

OpenStudy (mathmale):

I need just a sec.

OpenStudy (nicole143):

Okay

OpenStudy (mathmale):

Yes, yoiu're doing fine. But remember that Sqrt(4)=2 and Sqrt(36)=6. So your final answer becomes >>> ?

OpenStudy (nicole143):

8 over -4?

OpenStudy (mathmale):

Looks good! And 8 over -4 reduces to what? :)

OpenStudy (nicole143):

4 over -2 which goes to -2?

OpenStudy (mathmale):

Nice going, Nicole! I'm getting dizzy after having concentrated on OpenStudy questions for the past three hours or so, so forgive me if I leave soon. Looking forward to meeting you again online. All the best. MM

OpenStudy (nicole143):

Thank you for your help but I think we or I did something wrong, I don't have a simple -2 as an answer choice :( @mathmale

OpenStudy (mathmale):

Give me a mom. and I'll go thru the problem myself.

OpenStudy (nicole143):

Okay, these are my choice, maybe it will help:/

OpenStudy (nicole143):

*Choices

OpenStudy (mathmale):

Found the problem. We need to multiply as follows:\[(\sqrt{2}-\sqrt{6})(\sqrt{2}-\sqrt{6})\]

OpenStudy (nicole143):

\[\sqrt{4} - \sqrt{12} - \sqrt{12} - \sqrt{4} ?\]

OpenStudy (nicole143):

Oh, \[\sqrt{36}\]

OpenStudy (nicole143):

On the end not - 4 ha